# Matrix-free Second-order Optimization of Gaussian Splats with Residual Sampling

### v2

@@ 1 Introduction
Figure 1: We introduce a matrix-free second-order optimizer for Gaussian splatting. Notably, our dedicated optimizer converges significantly faster than Adam [adam] and already achieves reasonable renderings after very few seconds of training.
Photoreal novel view synthesis from multi-view images or video has attracted significant attention in recent years due to widely applicable downstream tasks in content creation, VR/XR, gaming, and the movie industry, to name a few. Here, Neural Radiance Fields (NeRF) [nerf] and 3D Gaussian Splatting (3DGS) [3dgs] mark a major milestone, due to their unprecedented quality leading to follow ups beyond view synthesis like VR rendering [VRNeRF], avatar creation [teotia2024gaussianheads], simultaneous localization and mapping (SLAM) [matsuki2024gaussianslam, adamkiewicz2022vision], and scene editing [bao2023sine, chen2024gaussianeditor].
However, NeRF-based models often require substantial training time, and researchers have developed various techniques to mitigate this problem. Some of them include neural hashing [instantngp], employing explicit scene modeling [plenoxels], improved sampling strategies [efficient_nerf], and tensor factorization methods [chen2022tensorf]. 3DGS [3dgs] instead does not rely on coordinate-based representations, but leverages a set of 3D Gaussians, which can be effectively rendered into image space using tile-based rasterization. Nonetheless, optimizing the parameters of each Gaussian can still take hours. Previous work focused on finding better densification strategies [taming, mcmc_splatting], quantization and compression of Gaussians [girish2024eagles, navaneet2023compact3d], and more efficient implementations for the backward pass [taming], which led to substantially reduced training times. However, all of them mostly rely on a first-order optimization routine, i.e. gradient descent or Adam [adam], and do not explore other (potentially second-order) alternatives.
@@
To this end, we formulate fitting the Gaussian parameters to the multi-view images as a non-linear least squares optimization problem and leverage the Levenberg-Marquardt (LM) algorithm for solving it. Concurrent work 3DGS-LM [hollein20243dgs] also adopts the LM optimizer with Gaussian splatting; however, it still suffers from the high storage and computation requirements of the Jacobian matrix. In addition, 3DGS-LM has to run the Adam [adam] optimizer first to provide fast convergence, since solving the numerical system arising from the normal equation is computationally expensive.
@@
Overall, with these improvements, our proposed second-order optimizer is both memory and compute efficient compared to 3DGS-LM [hollein20243dgs], suggesting that second-order optimization for 3DGS is a highly promising direction for further study. Our optimizer demonstrates improvements in settings with a low and moderate number of Gaussians over first-order optimizers as well. In summary, we propose a novel second-order optimizer for 3DGS with the following features:
@@ 2 Related Work
3D Scene Representation. Neural Radiance Fields (NeRFs) [nerf] represent the 3D geometry of a scene implicitly using a coordinate-based neural network. Subsequently, several variations of NeRFs have been proposed to improve visual quality [barron2021mipnerf] and to improve computational efficiency [sun2022direct, hedman2021baking, muller2022instant]. However, despite the significant advancements, NeRFs still face limitations, particularly in terms of their training and rendering efficiency. More recently, 3D Gaussian Splatting [3dgs] has been proposed, which employs a fully explicit representation, unlocking significant enhancements in rendering speed while maintaining similar qualitative results to NeRF-based methods. Recent advancements have further improved the rendering quality of 3DGS [lu2024scaffold, yu2024mip], or its efficiency [fan2025lightgaussian, lee2024compact, fang2024mini, rota2024revising, taming]. Notably, LightGaussian [fan2025lightgaussian] and C3DGS [lee2024compact] focus on pruning Gaussians, directly leading to improvements in rendering speed. Another related approach is to constrain the densification process, efficiently modeling the scene with less Gaussians, such as in Mini-Splatting [fang2024mini] and 3DGS-MCMC [mcmc_splatting]. Some other works aim to compress or quantize Gaussians [girish2024eagles, navaneet2023compact3d]. Taming3DGS [taming] steers the densification process in a controlled manner, while also devising a parallelized implementation for more efficient backpropagation in CUDA. Orthogonally to these works, we explore second-order optimizers for 3DGS optimization, which is challenging due to the high memory and computational demands associated with storing and inverting large Jacobian matrices. To achieve this, we exploit the sparsity of the Jacobian matrices of 3DGS.
@@
First-order Optimizers. Stochastic Gradient Descent (SGD) [sgd] was developed as a general optimization method and has been widely used in modern multi-dimensional deep learning tasks. With the addition of momentum [sgd_momentum], SGD can escape from local optima, a crucial feature in stochastic settings. However, a key limitation is that SGD applies the same update to all dimensions, which makes it difficult to adapt to complex problems. To address this, researchers have developed various preconditioners for the gradient information. Adagrad [adagrad] accumulates the squares of the gradients and preconditions the gradient with the square root of the accumulated values. RMSprop [rmsprop] replaces accumulation with an exponentially moving average, which leads to more stable training. Adam [adam], arguably the most popular optimizer today, combines momentum with the moving average of second-order moments. It has also become the default optimizer for many 3D tasks, including NeRF and 3DGS, and provides a strong baseline. However, despite the efficacy of first-order optimizers, there is strong incentive to explore second-order optimizers due to their attractive properties, as discussed next.
@@
Second-order Optimizers. Second-order optimizers often approximate the inverse of the Hessian matrix to precondition the gradient, and can offer significant advantages over first-order methods. They require far fewer iterations because they can obtain the second-order approximation of the loss landscape into account [kfac, martens2010deep, martens2011learning]. Additionally, second-order methods are often not as sensitive to learning rate values, and can often estimate it locally through line search algorithms [armijo1966minimization, hollein20243dgs, more1994line], or trust region methods [conn2000trust, nocedal1999numerical].
@@
where $\mathbf{g}\in\mathbb{R}^{n}$ is the gradient, $\mathbf{H}\in\mathbb{R}^{n\times n}$ is the Hessian of the loss function and $\boldsymbol{\beta}\in\mathbb{R}^{n}$ are the parameters to be updated, and $n$ is the number of parameters. In a naive implementation, storing the Hessian requires $O(n^{2})$ space, and computing its inverse takes $O(n^{3})$ time, making it impractical for large-scale optimization tasks. Instead, a numerical approximation of the inverse can be obtained using iterative solvers like conjugate gradient (CG) [hestenes1952methods]. This method can also eliminate the need to store the Hessian matrix, as CG only requires the results of matrix-vector multiplications [martens2010deep, martens2011learning]. In the literature, this class of methods is referred to as Hessian-free or matrix-free optimization.
Specifically, within the class of second-order optimizers, the Gauss-Newton (GN) algorithm [lecun2002efficient] is sometimes adopted to approximate the Hessian $\mathbf{H}$ with $\mathbf{J^{\top}J}$, where $\mathbf{J}$ is the Jacobian matrix. This method effectively approximates the second-order derivative only using first-order derivatives. Notably, the Levenberg-Marquardt (LM) algorithm [more2006levenberg] is an extension over GN, interpolating between GN and gradient descent, resulting in more stable optimization. Recently, some works [lan20253dgs, hollein20243dgs] also explored second-order optimizers for 3DGS optimization. 3DGS2 [lan20253dgs] explores an algorithm based on Newton’s method, but makes it computationally tractable by limiting the second-order computations to a small locality. Concurrent work 3DGS-LM [hollein20243dgs] employs the LM optimizer with a dedicated cache for intermediate Jacobian entries for fast computation. However, caching Jacobians results in quadratic memory requirements, as it scales with respect to both pixels and parameters. In addition, 3DGS-LM heavily relies on Adam optimizer for performance since processing a large cache is also expensive.
@@
3 Background - 3D Gaussian Splatting
3DGS is a point-based representation that can reconstruct 3D scenes with high fidelity. The 3D scene is represented using a set of Gaussians, where each Gaussian has an optimizable mean $\mathbf{x}$ and covariance $\boldsymbol{\Sigma}$:
$G(\mathbf{x})=e^{-\frac{1}{2}\mathbf{x}^{T}\boldsymbol{\Sigma}^{-1}\mathbf{x}}$ (2)
@@ 3 Background - 3D Gaussian Splatting
To render a pixel, the Gaussians are first projected onto the 2D image plane, and sorted according to their depth. A pixel value $C_{i}$ is computed using $\alpha$-blending, which combines the color $c$ and per-pixel opacity $\alpha$ of the projected Gaussians:
$C_{i}=$ $\sum_{n\leq\mathcal{S}}c_{n}\cdot\alpha_{n}\cdot\prod_{m<n}(1-\alpha_{m})$ (4)
, $\text{, where}\quad\alpha_{n}=o_{n}\boldsymbol{\pi}_{\text{cam}}\left(G(\mathbf{n})\right)$
where $\mathcal{S}$ represents the total number of depth-sorted Gaussians projected onto pixel $i$, and each $\alpha_{n}$ is obtained by multiplying Gaussian opacity $o_{n}$ and its 2D projection (via camera projection $\boldsymbol{\pi}_{\text{cam}}$).
@@
where $\mathcal{L}_{f}(\cdot)$ is a loss function that quantifies image differences. In this paper, we experiment with mean squared error (MSE) and structural similarity (SSIM) [ssim] loss functions. Additionally, 3DGS models the view-dependent effects with spherical harmonics. In this work, however, we assume each scene has Lambertian surface and disable the corresponding spherical harmonics levels.
@@ 4 Method
We first introduce the LM optimizer in Sec. 4.1, which we adopt in this work. Then, we derive why a naive implementation is not feasible for Gaussian splatting. To resolve this issue, in Sec. 4.2, we discuss our matrix-free approach to adapt the LM optimizer. In detail, we propose a GPU-parallelized conjugate gradient solver for our second-order 3DGS optimizer, which circumvents explicit storage of the Jacobian matrix, and solves the matrix inverse iteratively. Next, in Sec. 4.3, we introduce a new view sampling strategy to effectively approximate the full normal equation, thereby providing reliable update step directions by integrating information from multiple views. In Sec. 4.4, we present our residual sampling, providing an approximate loss function, which results in significantly faster convergence. Lastly, in Sec. 4.5, we introduce a heuristic to automatically determine the learning rate, which eliminates the need for line search algorithms that are commonly used in conjunction with second-order optimizers.
@@ 4.1 Levenberg-Marquardt Optimizer for 3DGS
In practice, naively solving the normal equation is not feasible, especially in large-scale numerical systems. First, the Jacobians scale quadratically with both the number of Gaussians and the number of pixels, and storing the Jacobians explicitly in memory quickly becomes infeasible. For example, using $100$ images at a resolution of $800\times 800$ pixels as training data, along with 10 00010\,000 Gaussians to model the 3D scene, would result in Jacobians exceeding $100$ TB in size. Even assuming a sparse storage format with a $99\%$ sparsity, the storage requirement would still exceed $1$ TB for a true LM optimizer. In addition, explicitly inverting the system matrix has O($P^{3}$) time complexity when implemented naively. Thus, computing the explicit inverse is often infeasible, and previous work has focused on solving the normal equation iteratively [martens2010deep, martens2011learning, byrd2011use], but they still remain computationally expensive, as discussed in the next section.
@@ 4.2 Matrix-free Preconditioned Conjugate Gradient (PCG) Solver for 3DGS
We first discuss our matrix-free PCG solver to adapt the LM optimizer. In this subsection, we present a GPU-parallelized conjugate gradient solver, which does not need to store the Jacobian matrix explicitly, avoiding the aforementioned memory issues. To achieve this, we solve Eq. 8 via the preconditioned conjugate gradient algorithm [hestenes1952methods], which only needs results of Jacobian-vector products, and finds the solution $\Delta\boldsymbol{\beta}$ iteratively.
To improve the condition number of the linear system in Eq. 8, we used the Jacobi preconditioner due to its simplicity and effectiveness. More specifically, we use $\frac{1}{\operatorname{diag}(\mathbf{J^{\top}J}+\lambda\mathbf{I})}\in\mathbb{R}^{P}$, as the precondition. We provide the implementation of the solver in CUDA, together with efficient Jacobian-vector product kernel, which is implemented with forward mode differentiation and dual numbers [baydin2018automatic, wengert1964simple]. See the supplementary material for details about the kernel design. The pseudocode of the optimizer is given in Alg. 1.
Although this optimizer is able to converge to the final solution in a limited number of steps, it is $4\times$ slower than our final method, as shown in Tab. 3. The main reason for this is that the conjugate gradient algorithm needs to run several iterations, therefore, we need to repeatedly compute the $\mathbf{J}^{\top}\mathbf{J}\mathbf{p}$ product arising on line $10$ of Alg. 1. This kind of computational complexity commonly arises when a second-order optimizer is used in large-scale numerical systems, and usually, an approximation method is used for efficiency. The common approximation methods are diagonal [becker1989improving, sophia, adahessian, pesky] and block-diagonal [kfac, roux2007topmoumoute, botev2017practical] approximation of the Hessian or $\mathbf{J^{\top}J}$ matrix.
@@ 4.3 View Sampling
Second-order methods like LM are typically used in deterministic settings where the full objective is evaluated at every iteration. If this is not the case, estimating local curvature can be problematic and become unreliable [byrd2011use]. This poses a challenge in the 3DGS setting, where the number of views can exceed hundreds, as incorporating all of them at the same time is infeasible. Yet, to compute meaningful gradients, we must find an effective way to approximate the full normal equation in Eq. 8.
@@
$f_{i}=[\tilde{x}_{i},\tilde{y}_{i},\tilde{z}_{i},\;{dx}_{i},{dy}_{i},{dz}_{i}]$ (10)
[…]

### v3

@@ 1 Introduction
Figure 1: We introduce a matrix-free second-order optimizer for Gaussian Splatting. Notably, our dedicated optimizer converges significantly faster than Adam [24] and already achieves reasonable renderings after very few seconds of training.
Photoreal novel view synthesis from multi-view images or video has attracted significant attention in recent years due to widely applicable downstream tasks in content creation, VR/XR, gaming, and the movie industry, to name a few. Here, Neural Radiance Fields (NeRF) [36] and 3D Gaussian Splatting (3DGS) [22] mark a major milestone, due to their unprecedented quality leading to follow ups beyond view synthesis like VR rendering [54], avatar creation [50], simultaneous localization and mapping (SLAM) [35, 1], and scene editing [3, 11].
However, NeRF-based models often require substantial training time, and researchers have developed various techniques to mitigate this problem. Some of them include neural hashing [39], employing explicit scene modeling [47], improved sampling strategies [21], and tensor factorization methods [10]. 3DGS [22] instead does not rely on coordinate-based representations, but leverages a set of 3D Gaussians, which can be effectively rendered into image space using tile-based rasterization. Nonetheless, optimizing the parameters of each Gaussian can still take hours. Previous work focused on finding better densification strategies [31, 23], quantization and compression of Gaussians [16, 41], and more efficient implementations for the backward pass [31], which led to substantially reduced training times. However, all of them mostly rely on a first-order optimization routine, i.e. gradient descent or Adam [24], and do not explore other (potentially second-order) alternatives.
@@
To this end, we formulate fitting the Gaussian parameters to the multi-view images as a non-linear least squares optimization problem and leverage the Levenberg-Marquardt (LM) algorithm for solving it. Concurrent work 3DGS-LM [20] also adopts the LM optimizer with Gaussian Splatting; however, it still suffers from the high storage and computation requirements of the Jacobian matrix. In addition, 3DGS-LM has to run the Adam [24] optimizer first to provide fast convergence, since solving the numerical system arising from the normal equation is computationally expensive.
@@
Overall, with these improvements, our proposed second-order optimizer is both memory and compute efficient compared to 3DGS-LM [20], suggesting that second-order optimization for 3DGS is a highly promising direction for further study. Our optimizer demonstrates improvements in settings with a low and moderate number of Gaussians over first-order optimizers as well. In summary, we propose a novel second-order optimizer for 3DGS with the following features:
@@ 2 Related Work
3D Scene Representation. Neural Radiance Fields (NeRFs) [36] represent the 3D geometry of a scene implicitly using a coordinate-based neural network. Subsequently, several variations of NeRFs have been proposed to improve visual quality [4] and to improve computational efficiency [49, 18, 40]. However, despite the significant advancements, NeRFs still face limitations, particularly in terms of their training and rendering efficiency. More recently, 3D Gaussian Splatting (3DGS) [22] has been proposed, which employs a fully explicit representation, unlocking significant enhancements in rendering speed while maintaining better qualitative results than NeRF-based methods. Recent advancements have further improved the rendering quality of 3DGS [30, 56], or its efficiency [14, 28, 15, 44, 31]. Notably, LightGaussian [14] and C3DGS [28] focus on pruning Gaussians, directly leading to improvements in rendering speed. Another related approach is to constrain the densification process, efficiently modeling the scene with less Gaussians, such as in Mini-Splatting [15] and 3DGS-MCMC [23]. Some other works aim to compress or quantize Gaussians [16, 41]. Taming3DGS [31] steers the densification process in a controlled manner, while also devising a parallelized implementation for more efficient backpropagation in CUDA. Orthogonally to these works, we explore second-order optimizers for 3DGS optimization, which is challenging due to the high memory and computational demands associated with storing and inverting large Jacobian matrices. To achieve this, we exploit the sparsity of the Jacobian matrices of 3DGS.
@@
First-order Optimizers. Stochastic Gradient Descent (SGD) [43] was developed as a general optimization method and has been widely used in modern multi-dimensional deep learning tasks. With the addition of momentum [46], SGD can overcome local optima, a crucial feature in stochastic settings. However, a key limitation is that SGD applies the same update to all dimensions, which makes it difficult to adapt to complex problems. To address this, researchers have developed various preconditioners for the gradient information. Adagrad [13] accumulates the squares of the gradients and preconditions the gradient with the square root of the accumulated values. RMSprop [51] replaces accumulation with an exponentially moving average, which leads to more stable training. Adam [24], arguably the most popular optimizer today, combines momentum with the moving average of second-order moments. It has also become the default optimizer for many 3D tasks, including NeRF and 3DGS, and provides a strong baseline. However, despite the efficacy of first-order optimizers, there is strong incentive to explore second-order optimizers due to their attractive properties, as discussed next.
@@
Second-order Optimizers. Second-order optimizers often approximate the inverse of the Hessian matrix to precondition the gradient, and can offer significant advantages over first-order methods. They require far fewer iterations because they can obtain the second-order approximation of the loss landscape into account [32, 34, 33]. Additionally, second-order methods are often not as sensitive to learning rate values, and can often estimate it locally through line search algorithms [2, 20, 37], or trust region methods [12, 42].
@@
where $\mathbf{g}\in\mathbb{R}^{n}$ is the gradient, $\mathbf{H}\in\mathbb{R}^{n\times n}$ is the Hessian of the loss function and $\boldsymbol{\beta}\in\mathbb{R}^{n}$ are the parameters to be updated, and $n$ is the number of parameters. In a naive implementation, storing the Hessian requires $O(n^{2})$ space, and computing its inverse takes $O(n^{3})$ time, making it impractical for large-scale optimization tasks. Instead, a numerical approximation of the inverse can be obtained using iterative solvers like conjugate gradient (CG) [19]. This method can also eliminate the need to store the Hessian matrix, as CG only requires the results of matrix-vector multiplications [34, 33]. In the literature, this class of methods is referred to as Hessian-free or matrix-free optimization.
Specifically, within the class of second-order optimizers, the Gauss-Newton (GN) algorithm [27] is sometimes adopted to approximate the Hessian $\mathbf{H}$ with $\mathbf{J^{\top}J}$, where $\mathbf{J}$ is the Jacobian matrix. This method effectively approximates the second-order derivative only using first-order derivatives. Notably, the Levenberg-Marquardt (LM) algorithm [38] is an extension over GN, interpolating between GN and gradient descent, resulting in more stable optimization. Recently, some works [26, 20] also explored second-order optimizers for 3DGS optimization. 3DGS2 [26] explores an algorithm based on Newton’s method, but makes it computationally tractable by limiting the second-order computations to a small locality. Concurrent work 3DGS-LM [20] employs the LM optimizer with a dedicated cache for intermediate Jacobian entries for fast computation. However, caching Jacobians results in quadratic memory requirements, as it scales with respect to both pixels and parameters. In addition, 3DGS-LM heavily relies on Adam optimizer for performance since processing a large cache is also expensive.
@@
3 Background - 3D Gaussian Splatting (3DGS)
3DGS is a point-based representation that can reconstruct 3D scenes with high fidelity. The 3D scene is represented using a set of Gaussians, where each Gaussian has an optimizable mean $\mathcal{X}$ and covariance $\boldsymbol{\Sigma}$:
$G(\mathbf{x})=e^{-\frac{1}{2}\left(\mathbf{x}-\mathcal{X}\right)^{T}\boldsymbol{\Sigma}^{-1}\left(\mathbf{x}-\mathcal{X}\right)}$ (2)
@@ 3 Background - 3D Gaussian Splatting (3DGS)
To render a pixel, the Gaussians are first projected onto the 2D image plane, and sorted according to their depth. A pixel value $\hat{I_{i}}$ is computed using $\alpha$-blending, which combines the color $c$ and per-pixel opacity $\alpha$ of the projected Gaussians:
$\hat{I_{i}}$= $\sum_{n\leq\mathcal{S}}c_{n}\cdot\alpha_{n}\cdot\prod_{m<n}(1-\alpha_{m})$ (4)
, $\text{, where}\quad\alpha_{n}=o_{n}\boldsymbol{\pi}_{\text{cam}}\left(G_{n}\right)$
where $\mathcal{S}$ represents the total number of depth-sorted Gaussians projected onto pixel $i$, and opacity contribution $\alpha_{n}$ is obtained by multiplying Gaussian base opacity $o_{n}$ with its 2D projection (via camera projection $\boldsymbol{\pi}_{\text{cam}}$).
@@
where $\mathcal{L}_{f}(\cdot)$ is a loss function that quantifies image differences. In this paper, we experiment with mean squared error (MSE) and structural similarity (SSIM) [52] loss functions. Additionally, 3DGS models the view-dependent effects with spherical harmonics. In this work, however, we assume each scene has Lambertian surface and disable the corresponding spherical harmonics levels.
@@ 4 Method
We first introduce the LM optimizer in Sec. 4.1, which we adopt in this work. Then, we derive why a naive implementation is not feasible for Gaussian Splatting. To resolve this issue, in Sec. 4.2, we discuss our matrix-free approach to adapt the LM optimizer. In detail, we propose a GPU-parallelized conjugate gradient solver for our second-order 3DGS optimizer, which circumvents explicit storage of the Jacobian matrix, and solves the matrix inverse iteratively. Next, in Sec. 4.3, we introduce a new view sampling strategy to effectively approximate the full normal equation, thereby providing reliable update step directions by integrating information from multiple views. In Sec. 4.4, we present our residual sampling, providing an approximate loss function, which results in significantly faster convergence. Lastly, in Sec. 4.5, we introduce a heuristic to automatically determine the learning rate, which eliminates the need for line search algorithms that are commonly used in conjunction with second-order optimizers. The overview of all the components is given in Fig. 2.
@@ 4.1 Levenberg-Marquardt Optimizer for 3DGS
In practice, naively solving the normal equation is not feasible, especially in large-scale numerical systems. First, the Jacobians scale quadratically with both the number of Gaussians and the number of pixels, and storing the Jacobians explicitly in memory quickly becomes infeasible. For example, using $100$ images at a resolution of $800\times 800$ pixels as training data, along with 10 00010\,000 Gaussians to model the 3D scene, would result in Jacobians exceeding $100$ TB in size. Even assuming a sparse storage format with a $99\%$ sparsity, the storage requirement would still exceed $1$ TB for a true LM optimizer. In addition, explicitly inverting the system matrix has O($P^{3}$) time complexity when implemented naively. Thus, computing the explicit inverse is often infeasible, and previous work has focused on solving the normal equation iteratively [34, 33, 9], but they still remain computationally expensive, as discussed in the next section.
@@ 4.2 Matrix-free Preconditioned Conjugate Gradient (PCG) Solver for 3DGS
We first discuss our matrix-free PCG solver to adapt the LM optimizer. In this subsection, we present a GPU-parallelized conjugate gradient solver, which does not need to store the Jacobian matrix explicitly, avoiding the memory issues mentioned in Sec. 4.1. To achieve this, we solve Eq. 8 via the preconditioned conjugate gradient algorithm [19], which only needs results of Jacobian-vector products, and finds the solution $\Delta\boldsymbol{\beta}$ iteratively.
To improve the condition number of the linear system in Eq. 8, we used the Jacobi preconditioner due to its simplicity and effectiveness. More specifically, we use $\frac{1}{\operatorname{diag}(\mathbf{J^{\top}J}+\lambda\mathbf{I})}\in\mathbb{R}^{P}$, as the precondition. We provide the implementation of the solver in CUDA, together with efficient Jacobian-vector product kernel, which is implemented with forward mode differentiation and dual numbers [5, 53]. See the supplementary material for details about the kernel design. The pseudocode of the optimizer is given in Alg. 1.
Although this optimizer is able to converge to the final solution in a limited number of steps, it is $4\times$ slower than our final method, as shown in Tab. 3. The main reason for this is that the conjugate gradient algorithm needs to run several iterations, therefore, we need to repeatedly compute the $\mathbf{J}^{\top}\mathbf{J}\mathbf{p}$ product arising on line $10$ of Alg. 1. This kind of computational complexity commonly arises when a second-order optimizer is used in large-scale numerical systems, and usually, an approximation method is used for efficiency. The common approximation methods are diagonal [6, 29, 55, 48] and block-diagonal [32, 45, 7] approximation of the Hessian or $\mathbf{J^{\top}J}$ matrix.
@@ 4.3 View Sampling
Second-order methods like LM are typically used in deterministic settings where the full objective is evaluated at every iteration. If this is not the case, estimating local curvature can be problematic and become unreliable [9]. This poses a challenge in the 3DGS setting, where the number of views can exceed hundreds, as incorporating all of them at the same time is infeasible. Yet, to compute meaningful gradients, we must find an effective way to approximate the full normal equation in Eq. 8.
@@
$f_{i}=[x_{i},y_{i},z_{i},\;{dx}_{i},{dy}_{i},{dz}_{i}]$ (10)
where $(x,y,z)$ is the cameras’ normalized location and (dx,dy,dz)({dx},{dy},{dz}) is the viewing direction. We then run K-Means clustering on $\{f_{i}\}_{i=1}^{N}$ to partition the cameras into batch size number of clusters (e.g., 8 clusters created for a batch size of 8). A camera is then randomly selected from every cluster and collected in the set $\mathcal{C}_{B}$, which serves as training input. This approach ensures that each batch captures a more balanced and diverse set of views, resulting in a more accurate estimation of the curvature information. As evidenced by Tab. 2, this method converges to higher scores compared to the random sampling of the cameras.